# LLM with Relation Classifier for Document-Level Relation Extraction

### v1

@@ Abstract
Large language models (LLMs) create a new paradigm for natural language processing. Despite their advancement, LLM-based methods still lag behind traditional approaches in document-level relation extraction (DocRE), a critical task for understanding complex entity relations. This paper investigates the causes of this performance gap, identifying the dispersion of attention by LLMs due to entity pairs without relations as a primary factor. We then introduce a novel classifier-LLM approach to DocRE. The proposed approach begins with a classifier specifically designed to select entity pair candidates exhibiting potential relations and thereby feeds them to LLM for the final relation extraction. This method ensures that during inference, the LLM’s focus is directed primarily at entity pairs with relations. Experiments on DocRE benchmarks reveal that our method significantly outperforms recent LLM-based DocRE models and achieves competitive performance with several leading traditional DocRE models.
@@ 1 Introduction
Document-level Relation Extraction (DocRE) stands as a critical yet formidable undertaking within the expansive landscape of Natural Language Processing (NLP). It delves into the intricate world of identifying and extracting semantic relationships that exist between entity pairs across an entire document, rather than confining the analysis to individual sentences. This task poses significant challenges due to the complex interplay of context, coreference resolution, and the long-range dependencies that are often embedded within the textual corpus. Previous DocRE models have relied on sophisticated neural network architectures to mimic human-like processes of comprehending and reasoning, including attention networks Zhang and Cheng (2022); Tan et al. (2022a), GNNs Li et al. (2020); Christopoulou et al. (2019), and have attained state-of-the-art performance Ma et al. (2023).
@@
Recently, researchers Sun et al. (2024) have delved into leveraging novel strategies to further boost the capabilities of Large Language Models (LLMs) in tackling DocRE. However, the LLM-based method still lags behind traditional approaches. Our meticulous observation highlights a pivotal issue: the DocRE task necessitates the construction of an exhaustive set of potential entity pairs, which we refer to as candidate space. Within this vast landscape of entity pairs, only a minute fraction actually express relations, creating a significant imbalance within the candidate space. As shown in Figure 1, our preliminary investigations, conducted on two widely adopted benchmark datasets, DocRED Yao et al. (2019) and Re-DocRED Tan et al. (2022b), underscored this phenomenon. The overwhelming majority of entity pairs labeled as NA (no relation), dominate the candidate space, potentially distracting LLMs from the minority that do encode meaningful relations. This skew in the data distribution is consequently recognized as a key contributor to the performance shortcomings exhibited by LLMs on DocRE.
Based on this finding, this paper introduces an innovative method LMRC, aimed at bridging the performance gap between the LLM-based DocRE methods and their conventional counterparts. Specifically, LMRC conceptualizes DocRE as a workflow comprising two key stages: Relation Candidate Proposal (RCP) and Relation Classification (RC). In the RCP stage, a pre-processing classifier is employed, which harnesses attention mechanism to filter out NA entity pairs that do not harbour any relation, significantly narrowing down the scope of candidate space. Subsequently, the RC stage leverages the power of LLMs to perform multi-classification within this refined candidate space, ensuring a more focused and accurate relation extraction process. Experimental results on the DocRED and Re-DocRED benchmarks showed that the proposed LMRC gains significant improvement over other LLM-based DocRE methods, suggesting its viability as a strategy for future DocRE.
Our core contributions are as follows:
  • Through a detailed analysis of the two benchmark datasets, DocRED and Re-DocRED, we identified that the main reason why LLM-based methods struggle to perform well on the DocRE task lies in the extreme imbalance of the candidate space.
  • We proposed LMRC approach, which first utilizes a pre-classifier to select potential relation-expressing entity pairs, and then leverages the reasoning capabilities of LLMs to complete relation classification.
  • Empirical experiments conducted on DocRED and Re-DocRED datasets have validated the feasibility and effectiveness of our method.
@@ 2 Related Works
Standard NLP methods for identifying relations in documents have included CNN and LSTMYao et al. (2019); Gu et al. (2017); Mandya et al. (2018), which are built upon the foundations laid by early deep neural networks. Jia et al. (2019) advocated for an entity-centric, multiscale approach, enabling a singular relation prediction for each entity tuple. More recently, approaches for document-level relation extraction have evolved around two main perspectives: Graph-based and Transformer-based. The concept of document graph was initially introduced by Quirk and Poon (2016). In this paradigm, words are represented as nodes, while linguistic relations (e.g., adjacency, syntactic dependencies) are captured as edges. Graph-based methods Zeng et al. (2020); Zhang et al. (2020) particularly excels in handling long-distant dependencies that often proved crucial in comprehending complex relations. In addition to modifying the graph structure, Xu et al. (2022) proposed a sentence importance estimation and focusing framework, which encourages models to focus on evidence sentences. In early Transformer-based methods, Verga et al. (2018) introduced Bi-affine Relation Attention Network to identify the relations between entities. Han and Wang (2020) utilized a document-level entity mask method augmented with type information to provide a more richer contextual understanding of the entities. Zhou et al. (2021) and Tan et al. (2022a) further leveraged attention mechanism to enhance entity representations, improving the accuracy of relation classification. Recently, Ma et al. (2023) extended previous research by integrating knowledge distillation strategy and achieved SOTA performance.
@@
Relation Extraction with LLMs   LLMs have demonstrated immense potential across multiple domains Jiang et al. (2024); Jing et al. (2024). Several recently proposed RE approaches have advocated the utilization of LLMs to output string encodings, thereby addressing the task in an interpretable manner. Paolini et al. (2021) introduced a framework that reframes many structured prediction tasks, including relation extraction, as a sequence-to-sequence problem. Wan et al. (2023) pushed the boundaries by prompting GPT-3 for relation extraction on datasets such as TACRED Zhang et al. (2017) and SciERC Luan et al. (2018), achieving SOTA performance. Wadhwa et al. (2023) extended this line of work by leveraging few-shot in-context learning within LLMs for sentence-level relation extraction.
@@
Previously, the application of LLMs to document-level relation extraction has been limited. Ozyurt et al. (2023) introduced a method for in-context few-shot relation extraction leveraging pre-trained language models, providing a viable paradigm for tackling DocRE with LLMs. Sun et al. (2024) attempted to employ chain-of-retrieval and denoising strategy to steer LLMs towards accurately identifying relations. During their research on GPT-3’s contextual learning abilities for document-level biomedical information extraction tasks, Gutierrez et al. (2022) uncovered a general issue in in-context learning when dealing with the NA category. Inspire by this discovery, we first focus on selecting potential relation-expressing entity pairs and then utilize the reasoning capability of LLMs to complete relation classification. As far as we know, our work is among the first to apply LLMs for document-level relation extraction.
@@
In Context Learning   In-context learning Brown et al. (2020) has emerged as a competitor against supervised baselines across a wide range of tasks. This approach circumvents the computational burden of weight updates by leveraging a handful of pertinent examples presented to the model. Recent works on in-context learning have mainly focused on optimizing prompt retrieval Liu et al. (2022); Rubin et al. (2022) and prompt ordering Lu et al. (2022). Typically, Wan et al. (2023) and Wadhwa et al. (2023) have evaluated the in-context learning performance of GPT-3 on sentence-level RE tasks. We conduct our experiment in 5.6 following their works.
@@ 3.2 Pre-Experiments
To investigate the reasons behind the under-performance of current LLMs in DocRE, we fine-tune LLaMA2-13B-Chat for this task and report the outcomes and key insights derived from applying this approach to the DocRED and Re-DocRED datasets.
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exists relations, select relations for the entity pairs from the relation set; if there is no relation, return None.
The format of the input entity pair is ‘(head entity| -| tail entity)’.
Your output format is ‘(head entity| relation/None| tail entity)’.
### Relation set:
{predefined relation set}
### Text:
{text}
### {number of entity pairs} Entity pairs:
{entity pairs}
### Response:
Table 1: Prompt for document-level relation extraction
Fine-tuning LLaMA2 In order to construct effective prompts for this task, we employ the following instructional guideline: Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exist relations, select relations for the entity pairs from the relation set; if there is no relation, return None.. The instruction is accompanied by an input that includes a predefined relation set, the text corresponding to the document, and the entity pairs that need to be classified. To prevent ambiguities and optimize token usage in our process, we use None as a substitute for NA and require the model to explicitly label entity pairs that do not contain relations as None. The complete prompt format is provided in Table 1.
Each document in DocRE task involves a considerable number of tokens, frequently exceeding the maximum limit supported by typical processing unit. To tackle this challenge, we devise a method to conduct relation extraction for each document $D$ via $\frac{n\times(n-1)}{k}$ inputs, where $n$ denotes the number of entities within document $D$, and the variable $k$ represents the capacity limit for entity pairs per input instance. We then integrate all entity pairs into the inputs adhering to the prescribed scheme to perform LoRA Hu et al. (2022) fine-tuning and testing [^1] Entities in the triplets returned by LLaMA2-13B-Chat are aligned to the dataset using thefuzz, and the relations generated not in the predefined relation set are considered incorrect. on LLaMA2-13B-Chat.
Results Statistics of DocRED and Re-DocRED are shown in Table 2. NA entity pairs, which do not exhibit relations, constitute a significant proportion in both datasets, leading to an imbalance in the candidate space. Further to the empirical observations by Lilong et al. (2024), our study extends the analysis to examine the model’s outputs from a distribution perspective, supported by experiments, aiming to uncover the fundamental reasons behind the observed under-performance. As demonstrated in Table 3, the number of triples generated by LLaMA2-13B-Chat is markedly less than that annotated in the dataset. This discrepancy suggests that LLMs such as LLaMA2 tend to label relation-expressing entity pairs as NA, leading to the poor outcomes.
@@ 4.2 Relation Classification
After identifying potential relation-expressing entity pairs in the RCP stage, we employ the method described in Section 3, leveraging LLMs to undertake supervised fine-tuning for relation classification. This approach involves a subtle modification to the previous prompting scheme, where we streamline the process by eliminating the NA category and change certain expressions. These adjustments are intended to further sharpen the model’s concentration on the classification task, ensuring a more precise and targeted performance.
Additionally, the optimization achieved through the RCP stage yields a substantial reduction in the number of inputs per document, transitioning from a complexity of $\frac{n\times(n-1)}{k}$ to $\frac{n^{\prime}}{k}$, where $n^{\prime}$ denotes the number of entity pairs identified in the RCP stage. This reduction, driven by the elimination of NA instances, not only enhances computational efficiency but also shortens the inference time of LLMs. Details can be found in Table 4.
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
This is a relation classification task. we will provide entity pairs that require relation classification. Your task is to select relations for each entity pair from the given relation set based on the information in the text. There may be multiple relations between an entity pair.
The format of the input entity pair is ‘(head entity| -| tail entity)’.
Your output format is ‘(head entity| relation| tail entity)’.
### Relation set:
{predefined relation set}
### Text:
{text}
### {number of entity pairs} Entity pairs:
{entity pairs}
### Response:
Table 4: Prompt for relation classification

### v2

@@ Abstract
Large language models (LLMs) have created a new paradigm for natural language processing. Despite their advancement, LLM-based methods still lag behind traditional approaches in document-level relation extraction (DocRE), a critical task for understanding complex entity relations within long context. This paper investigates the causes of this performance gap, identifying the dispersion of attention by LLMs due to entity pairs without relations as a key factor. We then introduce a novel classifier-LLM approach to DocRE. Particularly, the proposed approach begins with a classifier designed to select entity pair candidates that exhibit potential relations and then feed them to LLM for final relation classification. This method ensures that the LLM’s attention is directed at relation-expressing entity pairs instead of those without relations during inference. Experiments on DocRE benchmarks reveal that our method significantly outperforms recent LLM-based DocRE models and narrows the performance gap with state-of-the-art BERT-based models. [^1] Our code is publicly available at: https://github.com/wisper12933/LMRC.
@@ 1 Introduction
Document-level relation extraction (DocRE) is a critical yet challenging task within the broad field of natural language processing (NLP). It involves identifying and extracting semantic relationships between entity pairs across an entire document, rather than confining the analysis to individual sentences. This task poses significant challenges due to the coreference resolution and long-distance dependencies inherent in the textual corpus. Traditional DocRE models have relied on sophisticated neural network architectures to simulate human-like comprehending and reasoning processes, such as GNNs Li et al. (2020); Christopoulou et al. (2019) and attention networks Zhang and Cheng (2022); Tan et al. (2022a), achieving state-of-the-art performance Ma et al. (2023). Recently, the robust semantic comprehension and generation capabilities of large language models (LLMs) have emerged as a promising avenue for advancing DocRE. Due to extensive pre-training, LLMs possess a broader coverage of knowledge across various fields and demonstrate enhanced versatility. With minimal fine-tuning, they can quickly adapt to domain-specific tasks. Besides, the textual output and in-context learning abilities of LLMs enable them to accomplish tasks in a more interpretable manner. Building on these advantages, researchers Lilong et al. (2024); Sun et al. (2024); Li et al. (2023) have intensively explored innovative strategies to further enhance the capabilities of LLMs in addressing complex DocRE tasks.
@@
Despite these efforts, the performance of LLM-based methods still fall short of traditional approaches. Our detailed analysis reveals a key issue: DocRE tasks necessitate the construction of an exhaustive set of potential entity pairs, which we refer to as candidate space. Within this large set of entity pairs, only a small subset expresses relations, leading to a significant imbalance in the candidate space. As shown in Figure 1, our preliminary investigations, conducted on two widely adopted benchmark datasets, DocRED Yao et al. (2019) and Re-DocRED Tan et al. (2022b), underscores this phenomenon. The overwhelming majority of entity pairs labeled as NA (no relation), dominates the candidate space, potentially distracting LLMs from the minority that do encode meaningful relations. Consequently, this data distribution skew is recognized as a major factor contributing to the performance shortcomings exhibited by LLMs on DocRE.
Based on this finding, this paper introduces a novel method named LMRC (LLM with Relation Classifier), which aims at bridging the performance gap between LLM-based DocRE methods and their conventional counterparts. Specifically, LMRC conceptualizes DocRE as a workflow comprising two stages: Relation Candidate Proposal (RCP) and Relation Classification (RC). In the RCP stage, a pre-processing classifier is applied, which explicitly harnesses attention mechanism to filter out NA entity pairs that do not harbour any relation. This strategic filtering not only drastically reduces the size of candidate space but also ensures that the subsequent stage operates on a more relevant set of entity pairs. Subsequently, the RC stage leverages the remarkable capabilities of LLMs to perform multi-classification within this refined candidate space, ensuring a more focused and accurate relation extraction process. Experimental results on the DocRED and Re-DocRED benchmarks shows that our proposed LMRC methodology gains significant improvement over other LLM-based DocRE methods, suggesting its viability as a strategy for future DocRE.
Our main work and contributions are as follows:
  • This study empirically found that the key reason why LLM-based methods struggle to perform well on DocRE tasks lies in the imbalance of the candidate space.
  • The proposed LMRC approach introduces a pre-classifier into LLMs to greatly narrow the performance gap with state-of-the-art BERT-based DocRE models.
  • Empirical experiments conducted on DocRED and Re-DocRED datasets showed the feasibility and effectiveness of The proposed LMRC approach.
@@ 2 Related Works
Document-level Relation Extraction Traditional NLP methods for identifying relations in documents, such as CNNs and LSTMs Yao et al. (2019); Gu et al. (2017); Mandya et al. (2018), are built on the foundations of early deep neural networks. Specifically, Jia et al. Jia et al. (2019) advocated for an entity-centric, multiscale approach, enabling a singular relation prediction for each entity tuple. More recently, approaches for document-level relation extraction have evolved around two main perspectives: Graph-based and Transformer-based. The concept of document graph was initially introduced by Quirk and Poon Quirk and Poon (2016). In this paradigm, words are represented as nodes, while linguistic relations (e.g., adjacency, syntactic dependencies) are captured as edges. Sun et al. Sun et al. (2022) proposed a method that combines dual-channel and hierarchical graph convolutional network, utilizing GCN to gather pertinent information distributed across the document to enhance the inference of implicit relations. Peng et al. Peng et al. (2022) introduced a subgraph reasoning framework that simplifies various paths into a streamlined subgraph structure, facilitating various reasoning processes. Specially, graph-based methods GAIN Zeng et al. (2020) and LSR Nan et al. (2020) excels in handling long-distant dependencies that often proved crucial in comprehending complex relations. In addition to modifying the graph structure, Xu et al. Xu et al. (2022) proposed a sentence importance estimation and focusing framework, which encourages models to focus on evidence sentences. Among early Transformer-based methods, Verga et al. Verga et al. (2018) introduced Bi-affine Relation Attention Network to identify the relations between entities. Han and Wang Han and Wang (2020) utilized a document-level entity mask method augmented with type information to provide a more richer contextual understanding of the entities. Yu et al. Yu et al. (2022) accounted for the importance of interactions among mentions and used a relation-specific mention attention network to incorporate relation into prototype representations. Zhou et al. Zhou et al. (2021) and Tan et al. Tan et al. (2022a) further leveraged attention mechanism to enhance entity representations, ultimately improving the accuracy of relation classification. Recently, Ma et al. Ma et al. (2023) extended previous research by integrating knowledge distillation strategy and achieved SOTA performance.
@@
Relation Extraction with LLMs   Several recently proposed RE approaches have advocated the utilization of conditional generative models to output string encodings, thereby addressing the task in an interpretable manner. Paolini et al. Paolini et al. (2021) introduced a framework that reframes many structured prediction tasks, including relation extraction, as a sequence-to-sequence problem. Wan et al. Wan et al. (2023) pushed the boundaries by prompting GPT-3 for relation extraction on datasets such as TACRED Zhang et al. (2017) and SciERC Luan et al. (2018). Wadhwa et al. Wadhwa et al. (2023) extended this line of work by leveraging few-shot in-context learning for sentence-level relation extraction. They also evaluated Flan-T5 to prove the potential of fine-tuning smaller scale models for accomplishing this challenging task.
@@
Previously, the application of LLMs to document-level relation extraction has been limited. Ozyurt et al. Ozyurt et al. (2023) introduced a method for in-context few-shot relation extraction leveraging pre-trained language models, providing a viable paradigm for tackling DocRE with LLMs. Xue et al. Lilong et al. (2024) guided LLMs to extract triplets from documents using a pipeline approach, starting with the extraction of head entity, followed by identifying the relation, and concluding with the retrieval of tail entity. Sun et al. Sun et al. (2024) attempted to employ chain-of-retrieval and denoising strategy to steer LLMs in understanding relations and generating high-quality synthetic data. During their research on GPT-3’s contextual learning abilities for document-level biomedical information extraction tasks, Gutierrez et al. Gutierrez et al. (2022) uncovered a general issue in in-context learning when dealing with the NA category. Inspired by this discovery, we first focus on selecting potential relation-expressing entity pairs and then utilize the reasoning capabilities of LLMs to complete relation classification. As far as we know, our work is among the first to apply LLMs for document-level relation extraction tasks.
@@
In-Context Learning   In-context learning Brown et al. (2020) has emerged as a competitor against supervised baselines across a wide range of tasks including semantic parsing, machine translation and question answering Liu et al. (2022); Zhao et al. (2021); Shin et al. (2021). This approach circumvents the computational burden of weight updates by leveraging a handful of pertinent examples presented to the model. Recent works on in-context learning have mainly focused on optimizing prompt design Perez et al. (2021), prompt ordering Lu et al. (2022) and prompt retrieval Liu et al. (2022); Rubin et al. (2022). Typically, Wan et al. Wan et al. (2023) and Wadhwa et al. Wadhwa et al. (2023) have conducted comprehensive evaluations of the in-context learning capabilities of GPT-3 on sentence-level RE tasks. We conduct our experiment in 5.6 following their pioneering works.
@@ 3.2 Pre-Experiments
To investigate the reasons behind the under-performance of current LLMs in DocRE, we directly fine-tune LLaMA2-13B-Chat and report the outcomes and key insights derived from applying this approach to DocRED and Re-DocRED datasets.
Fine-tuning LLaMA2 In order to construct effective prompts for this task, we employ the following instructional guideline: Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exist relations, select relations for the entity pairs from the relation set; if there is no relation, return None. The instruction is accompanied by an input that includes a predefined relation set, the text corresponding to the document, and the entity pairs that need to be classified. To prevent ambiguities and optimize token usage in our process, we use None as a substitute for NA and require the model to explicitly label entity pairs that do not contain relations as None. The complete prompt format is provided in Appendix A.
Each document in DocRE task involves a considerable number of tokens, frequently exceeding the maximum limit supported by typical processing unit. To tackle this challenge, we devise a method to conduct relation extraction for each document $D$ via $\frac{n\times(n-1)}{k}$ inputs, where $n$ denotes the number of entities within document $D$, and the variable $k$ represents the capacity limit for entity pairs per input instance. We then integrate all entity pairs into the inputs adhering to the prescribed scheme to perform LoRA Hu et al. (2022) fine-tuning and testing. In addition, entities of the triplets returned by models are aligned to the annotations in the dataset using thefuzz, and the relations generated not in the predefined relation set are considered incorrect.
Results   Statistics of DocRED and Re-DocRED are shown in Table 1. NA entity pairs, which do not exhibit relations, constitute a significant proportion in both datasets, leading to an imbalance in the candidate space. Further to the empirical observations by Lilong et al. (2024), our study extends the analysis to examine the model’s outputs from a distribution perspective, supported by experiments, aiming to uncover the fundamental reasons behind the observed under-performance. As demonstrated in Table 2, the number of triples generated by LLaMA2-13B-Chat is markedly less than that annotated in the dataset. This discrepancy suggests that LLMs tend to label relation-expressing entity pairs as NA, leading to poor outcomes.
@@ 4.2 Relation Classification
After identifying potential relation-expressing entity pairs in the RCP stage, we employ the method described in Section 3, leveraging LLMs to undertake supervised fine-tuning for relation classification. This approach involves a subtle modification to the previous prompting scheme, where we streamline the process by eliminating the NA category and change certain expressions. These adjustments are intended to further sharpen model’s concentration on the classification task, ensuring a more precise and targeted performance. Details can be found in Appendix A.
Additionally, the optimization achieved through the RCP stage yields a substantial reduction in the number of inputs per document, transitioning from a complexity of $\frac{n\times(n-1)}{k}$ to $\frac{n^{\prime}}{k}$, where $n^{\prime}$ denotes the number of entity pairs identified in the RCP stage. This reduction, driven by the elimination of NA instances, not only enhances computational efficiency but also shortens the inference time.